# Position: On-Premises LLM Deployment Demands a Middle Path: Preserving Privacy Without Sacrificing Model Confidentiality

### v2

@@ A.3 Proof of Theorem 1
$\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ $=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}\right]^{1/2}$ (22a)
$=\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{u}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (22b)
$\geq\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (22c)
$=\left[2-2\frac{\frac{\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p]\right|}}{\sqrt{1+\frac{\left|\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]\right|^{2}}{n\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])\right|^{2}}}}\right]^{1/2}$ (22d)
$=\left[2-2\frac{\frac{2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|}{\sqrt{n}(1+\beta_{D})^{L}\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|}}{\sqrt{1+\frac{2^{2[(1-\alpha)L-1]}}{n(1+\beta_{D})^{2L}}\frac{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|^{2}}{\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|^{2}}}}\right]^{1/2},$ (22e)
where equation (22b) is based on (18), equation (22e) is based on (21) and (14). When $\alpha>\log_{2}{\frac{2}{1+\beta_{D}}}$, we have $\lim_{L\to\infty}\left(\frac{2^{1-\alpha}}{1+\beta_{D}}\right)^{L}=0$. Thus we have $\lim_{L\to\infty}\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star}\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}$. This indicates that the $p$-th column of the output matrix $f(\mathbf{X}^{\star})$ is not parallel to $\mathbf{I}_{n}$ for any $p$. This further indicates that the output matrix does not have the identical vector in each row. ∎
@@ Appendix B Experiment Details
The foundation models we use in our experiments are selected from open-source repositories, and Table 3 shows the basic information of the models and their sources. Specifically, we employ Llama2-70B-chat [^1] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, Llama2-7B-chat [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and Mistral-7B-v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1. For smaller models, we select Phi-2444https://huggingface.co/microsoft/phi-2 and Phi-1.5555https://huggingface.co/microsoft/phi-1_5. We also consider OPT model [^6] https://huggingface.co/facebook/opt-350m, which has only 350 million parameters and 24 decoder layers.
@@
Model Size Decoder Layers
Llama2-70B-chat (Touvron et al., 2023) 70B 80
Llama2-7B-chat (Touvron et al., 2023) 7B 32
Mistral-7B-v0.1 (Jiang et al., 2023) 7B 32
Phi-2 (Abdin et al., 2024) 2.7B 32
Phi-1.5 (Li et al., 2023) 1.3B 24
OPT (Zhang et al., 2022) 350M 24
@@
Base 51k Distillation Dataset. We ensure dataset coverage and reliability by using a 1:1 ratio of the MMLU auxiliary training set  [^7] https://github.com/hendrycks/test and Alpaca dataset  [^8] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json, extracting 25.5k samples from each. From the MMLU auxiliary training data (Hendrycks et al., 2021), we sample 50%, and from Alpaca (Taori et al., 2023), we use a step size of 2 to enhance diversity. The datasets are then formatted for model training, applying Alpaca and MMLU prompts from Table 4.
Table 4: Prompts for Alpaca and MMLU auxiliary training data
@@
Validation Datasets. Table 5 outlines the composition of the validation dataset. We extracted 50% from each of the 57 MMLU validation sub-datasets, totaling 1.5K instances, paired with Alpaca data selected using a step size of 751.
Table 5: Composition of validation datasets of different sizes
@@
Total Length 1516 1516
@@ B.5 Evaluation Benchmarks
Evaluation on Customizabilities. We assess the customizability of models across six domains, as detailed in Table 6. Each domain includes specific benchmarks and metrics designed to evaluate different aspects of the model’s performance in relation to customizability. In particular, for evaluating medical capabilities, we select two subcategories from the MMLU benchmark that are related to the medical domain: mmlu_anatomy and mmlu_professional_medicine. For assessing legal reasoning, we select 10 multiple-choice and judgment-based subcategories from Legalbench. The performance of the model in these legal tasks is measured using perplexity, following the prompt structure provided by Legalbench. Specifically, the selected subcategories include: cuad_audit_rights, canada_tax_court_outcomes, definition_classification, cuad_affiliate_license-licensee, learned_hands_business, contract_nli_survival_of_obligations, contract_nli_explicit_identification, contract_nli_confidentiality_of_agreement, hearsay, and contract_qa.
@@
Code HumanEval Pass@1 0 (Chen et al., 2021)
MBPP Pass@1 1 (Austin et al., 2021)
Math GSM8K Exact Match 8 (Cobbe et al., 2021)
Medical MMLU_Medical Accuracy 5 (Hendrycks et al., 2021)
Finance FPB F1 0 (Wang et al., 2023a)
Law LegalBench Accuracy 0 (Guha et al., 2023)
Alignment MT-Bench Score (GPT-4) (Zheng et al., 2023)
Evaluation on Security. We follow the Llama-2 report (Touvron et al., 2023) to evaluate the distilled model, including 16 benchmarks, which are categorized into 6 groups. Table 7 summarizes the functionality benchmarks used in our experiments, along with their test methods and performance metrics. Our model ranks choices in multiple-choice tasks and generates answers for open-ended generation tasks.
@@
Domain Benchmark Metric n-shot Reference
Commonsense Reasoning PIQA Accuracy 0 (Bisk et al., 2020)
Hellaswag Accuracy 0 (Zellers et al., 2019)
Winogrande Accuracy 0 (Sakaguchi et al., 2019)
ARC_easy Accuracy 0 (Clark et al., 2018)
ARC_challenge Accuracy 0 (Clark et al., 2018)
Reading Comprehension OpenBookQ Accuracy 0 (Mihaylov et al., 2018)
LAMBADA Accuracy 0 (Paperno et al., 2016)
BoolQ Accuracy 0 (Clark et al., 2019)
SQuADv2 HasAns_EM 2 (Rajpurkar et al., 2018)
SQuADv2 HasAns_F1 2 (Rajpurkar et al., 2018)
World Knowledge NaturalQuestions Exact Match 5 (Kwiatkowski et al., 2019)
TriviaQA Exact Match 5 (Joshi et al., 2017)
Code HumanEval Pass@1 0 (Chen et al., 2021)
MBPP Pass@1 1 (Austin et al., 2021)
Math GSM8K Exact Match 8 (Cobbe et al., 2021)
General Ability MMLU Accuracy 5 (Hendrycks et al., 2021)
BBH Accuracy 3 (Suzgun et al., 2022)
@@ B.6 Model Customization
Datasets. To fine-tune the models for domain-specific tasks, we utilized several datasets tailored to different sectors, including Code (Zheng et al., 2024b), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023b), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). Table 8 lists the customization training datasets used in the experiments. For the code domain, we combine the datasets from CodeFeedback and CodeAlpaca. For law and finance, we merge all training datasets from Legalbench and FinGPT respectively. These datasets are then prepared for model training using the Alpaca prompts outlined in Table 4. Additionally, we randomly select 3,000 samples to serve as the validation dataset.
@@
Domain Dataset Name Size Reference
Code CodeFeedback 156k  (Zheng et al., 2024a)
CodeAlpaca 20k (Chaudhary, 2023)
Math MathInstruction 262K (Yue et al., 2023)
Medical MedMCQA 183k (Zhang et al., 2023)
Law Legalbench 90k (Guha et al., 2023)
Finance FinGPT 204k (Wang et al., 2023a)
Alignment Ultrafeedback 62k (Cui et al., 2024)
Customization Training Hyperparameters. In model customization, we use different hyperparameters depending on the model size. For LLaMA2-70B, we apply QLoRA with the settings outlined in Table 9, while for 7B models, we use LoRA. For smaller models like Phi2 and Phi-1.5, we fine-tune all model parameters. For LLaMA2-70B, we fine-tune it as a quantized 4-bit model over 1 epoch, starting with a learning rate of $1.5\times 10^{-6}$. For the 7B models, we train for 3 epochs, with a seed value of 42. The training setup includes a weight decay of 0.1, a batch size of 128, a warmup ratio of 0.03, and input sequences of 512 tokens, following standard experimental practices (Hu et al., 2021). For Phi2 and Phi-1.5, we use the training hyperparameters from the LLaMA2 report. We employ the AdamW optimizer with a cosine learning rate scheduler, starting with a learning rate of $2\times 10^{-5}$, a weight decay of 0.1, a batch size of 128, and use bfloat16 precision for 512-token input sequences. Specifically, for alignment, we follow SimPO (Meng et al., 2024) and set the preference parameters $\beta=2$ and $\gamma=1$. The learning rate is $1\times 10^{-6}$ for LLaMA2-70B and $5\times 10^{-7}$ for the 7B and smaller models. All experiments are conducted using the LLaMA-Factory on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, with PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Table 9: The Hyperparameters for Customization Training.
@@ B.8 Security Across Secure Sizes
50M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ $MLP$ 0
@@
We further computed $\Delta\text{ADR}$ by close-sourcing varying quantities and proportions of parameters under FT-all attacks on three additional models. As shown in Figure 8 and Figure LABEL:fig:appendix:appendix-specific-domain(a), we observed the same pattern as with Llama2-7B, where security emerges once a sufficient number of parameters are secured. For example, on Mistral-7B, security occurs after secure-sourcing 100 million parameters, which is less than a single decoder layer. Secure-sourcing fewer parameters leads to a notable drop in security, with $\Delta\text{ADR}$ rising to around 40%. Beyond this threshold, security stabilizes near 0% $\Delta\text{ADR}$. This pattern holds across all models, highlighting a critical threshold for effective secure-source. Furthermore, different architectures require varying secure-sourcing quantities to achieve security, even with similar model sizes. For instance, Mistral-7B reaches security by secure-sourcing 100 million parameters, Llama2-7B requires 200 million, and Phi-1.5 needs a higher rate of 7%, compared to 3% for Llama2-7B.
@@
We explore how secured parameter ratio impacts the model security in Llama2-7B, as shown in Figure LABEL:fig:appendix:appendix-specific-domain(b). For instance, technical skills such as Math show earlier transitions, with security emerging at 1% parameters secured, whereas domains such as Commonsense Reasoning require hiding 3%. In summary, secure-sourcing a small portion of parameters can provide sufficient security against model distillation, meanwhile, technical capabilities tend to be more challenging to distill than other domains.

### v3

@@ A.3 Proof of Theorem 1
$\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}$ (22a)
$=\left[2-\frac{2\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left\|f(\mathbf{X}^{\star})\left[p\right]\right\|_{2}}\right]^{1/2}$ (22b)
$=\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\cdot\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\begin{array}[]{l}\frac{1}{n}\left(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]\right)^{2}+\left(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p]\right)^{2}\\[2.15277pt] +\left(\bm{u}^{\star\top}f(\mathbf{X}^{\star})[p]\right)^{2}\end{array}}}\right]^{1/2}$ (22e)
$\geq\left[2-\frac{2\mathbb{I}_{n}^{\top}}{\sqrt{n}}\frac{f(\mathbf{X}^{\star})[p]}{\sqrt{\frac{1}{n}(\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p])^{2}+(\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])^{2}}}\right]^{1/2}$ (22f)
$=\left[2-2\frac{\frac{\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]}{\sqrt{n}\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p]\right|}}{\sqrt{1+\frac{\left|\mathbb{I}_{n}^{\top}f(\mathbf{X}^{\star})[p]\right|^{2}}{n\left|\bm{v}^{\star\top}f(\mathbf{X}^{\star})[p])\right|^{2}}}}\right]^{1/2}$ (22g)
$=\left[2-2\frac{\frac{2^{(1-\alpha)L-1}\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|}{\sqrt{n}(1+\beta_{D})^{L}\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|}}{\sqrt{1+\frac{2^{2[(1-\alpha)L-1]}}{n(1+\beta_{D})^{2L}}\frac{\left|\mathbb{I}_{n}^{\top}\hat{\varphi}_{\alpha L}\circ\varphi_{\alpha L-1}\circ\cdots\circ\varphi_{1}(\mathbf{X}^{\star})[p]\right|^{2}}{\left|\bm{v}^{\star\top}\mathbf{X}^{\star}[p]\right|^{2}}}}\right]^{1/2},$ (22h)
where equation (22e) is based on (18), equation (22h) is based on (21) and (14). When $\alpha>\log_{2}{\frac{2}{1+\beta_{D}}}$, we have $\lim_{L\to\infty}\left(\frac{2^{1-\alpha}}{1+\beta_{D}}\right)^{L}=0$. Thus we have $\lim_{L\to\infty}\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star}\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}$. This indicates that the $p$-th column of the output matrix $f(\mathbf{X}^{\star})$ is not parallel to $\mathbf{I}_{n}$ for any $p$. This further indicates that the output matrix does not have the identical vector in each row. ∎
@@ Appendix B Experiment Details
The foundation models we use in our experiments are selected from open-source repositories, and Table 4 shows the basic information of the models and their sources. Specifically, we employ Llama2-70B-chat [^1] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf (Touvron et al., 2023), Llama2-7B-chat [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023), and Mistral-7B-v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1 (Jiang et al., 2023). For smaller models, we select Phi-2444https://huggingface.co/microsoft/phi-2 (Abdin et al., 2024) and Phi-1.5555https://huggingface.co/microsoft/phi-1_5 (Li et al., 2023). We also consider OPT model [^6] https://huggingface.co/facebook/opt-350m (Zhang et al., 2022), which has only 350 million parameters and 24 decoder layers.
Model Size Decoder Layers
Llama2-70B-chat 70B 80
Llama2-7B-chat 7B 32
Mistral-7B-v0.1 7B 32
Phi-2 2.7B 32
Phi-1.5 1.3B 24
OPT 350M 24
@@
Base 51k Distillation Dataset. We ensure dataset coverage and reliability by using a 1:1 ratio of the MMLU  [^7] https://github.com/hendrycks/test (Hendrycks et al., 2021) auxiliary training set and Alpaca dataset  [^8] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json (Taori et al., 2023), extracting 25.5k samples from each. From the MMLU auxiliary training data, we sample 50%, and from Alpaca, we use a step size of 2 to enhance diversity. The datasets are then formatted for model training, applying Alpaca and MMLU prompts from Table 5.
@@
Table 5: Prompts for Alpaca and MMLU auxiliary training data
Extra Distillation Datasets. To enhance dataset diversity, the 100K, 200K, 300K, and 500K datasets integrate additional specialized sources. As detailed in Table 6, these sources include Baize (Xu et al., 2023) (158K English multi-turn conversations via ChatGPT’s self-chat), MathInstruct (Yue et al., 2023) (260K curated math instruction instances focusing on hybrid reasoning), and OpenOrca (Mukherjee et al., 2023) (augmented FLAN collection with 1M GPT-4 completions and 3.2M GPT-3.5 completions). These enrichments are intended to support complex computational and theoretical tasks, offering broader topic coverage.
Raw Data Set 51k 100k 200k 300k 500k
Alpaca 25.5 50 40 50 50
MMLU auxiliary training set 25.5 50 40 100 100
Baize-MedQuAD 0 0 40 50 50
Baize-Quora 0 0 40 50 50
Baize-Stackoverflow 0 0 40 50 50
MathInstruct 0 0 4 6 20
OpenOrca 0 0 0 0 180
Table 6: Composition of variously sized datasets
Validation Datasets. Table 7 outlines the composition of the validation datasets. For Validation Dataset 1, we extracted 50% from each of the 57 MMLU validation sub-datasets, totaling 1.5K instances, paired with Alpaca data selected using a step size of 751. This dataset is used with the 51K and 100K training sets. For larger training sets (200K, 300K, and 500K), Validation Dataset 2 was created by adding 400 instances from three Baize subsets, expanding the validation set to 4.0K.
@@
Baize-MedQuAD 0 850
Baize-Quora 0 850
Baize-Stackoverflow 0 850
Total Length 1516 4066
Table 7: Composition of validation datasets of different sizes
@@ B.5 Evaluation Benchmarks
Evaluation on Customizabilities. We assess the customizability of models across six domains, as detailed in Table 8. Each domain includes specific benchmarks and metrics designed to evaluate different aspects of the model’s performance in relation to customizability. In particular, for evaluating medical capabilities, we select two subcategories from the MMLU benchmark that are related to the medical domain: mmlu_anatomy and mmlu_professional_medicine. For assessing legal reasoning, we select 10 multiple-choice and judgment-based subcategories from Legalbench. The performance of the model in these legal tasks is measured using perplexity, following the prompt structure provided by Legalbench. Specifically, the selected subcategories include:
  • cuad_audit_rights
  • canada_tax_court_outcomes
  • definition_classification
  • cuad_affiliate_license-licensee
  • learned_hands_business
  • contract_nli_survival_of_obligations
  • contract_nli_explicit_identification
  • contract_nli_confidentiality_of_agreement
  • hearsay
  • contract_qa
Domain Benchmark Metric n-shot Reference
Code HumanEval Pass@1 0 Chen et al. (2021)
MBPP Pass@1 1 Austin et al. (2021)
Math GSM8K Exact Match 8 Cobbe et al. (2021)
Medical MMLU_Medical Accuracy 5 Hendrycks et al. (2021)
Finance FPB F1 0 Wang et al. (2023b)
Law LegalBench Accuracy 0 Guha et al. (2023)
Alignment MT-Bench Score (GPT-4) Zheng et al. (2023)
@@
Evaluation on Security. We follow the Llama-2 report Touvron et al. (2023) to evaluate the distilled model, including 16 benchmarks, which are categorized into 6 groups. Table 9 summarizes the functionality benchmarks used in our experiments, along with their test methods and performance metrics. Our model ranks choices in multiple-choice tasks and generates answers for open-ended generation tasks.
@@
Commonsense Reasoning PIQA Accuracy 0 Bisk et al. (2020)
Hellaswag Accuracy 0 Zellers et al. (2019)
Winogrande Accuracy 0 Sakaguchi et al. (2019)
ARC_easy Accuracy 0 Clark et al. (2018)
ARC_challenge Accuracy 0 Clark et al. (2018)
Reading Comprehension OpenBookQ Accuracy 0 Mihaylov et al. (2018)
LAMBADA Accuracy 0 Paperno et al. (2016)
BoolQ Accuracy 0 Clark et al. (2019)
SQuADv2 HasAns_EM 2 Rajpurkar et al. (2018)
SQuADv2 HasAns_F1 2 Rajpurkar et al. (2018)
World Knowledge NaturalQuestions Exact Match 5 Kwiatkowski et al. (2019)
TriviaQA Exact Match 5 Joshi et al. (2017)
Code HumanEval Pass@1 0 Chen et al. (2021)
MBPP Pass@1 1 Austin et al. (2021)
Math GSM8K Exact Match 8 Cobbe et al. (2021)
General Ability MMLU Accuracy 5 Hendrycks et al. (2021)
BBH Accuracy 3 Suzgun et al. (2022)
@@ B.6 Model Customization
Datasets. To fine-tune the models for domain-specific tasks, we utilized several datasets tailored to different sectors, including Code (Zheng et al., 2024a), Math (Yue et al., 2023), Medical (Zhang et al., 2023), Finance (Wang et al., 2023a), Law (Guha et al., 2024), and Alignment (Meng et al., 2024). Table 10 lists the customization training datasets used in the experiments. For the code domain, we combine the datasets from CodeFeedback and CodeAlpaca. For law and finance, we merge all training datasets from Legalbench and FinGPT respectively. These datasets are then prepared for model training using the Alpaca prompts outlined in Table 5. Additionally, we randomly select 3,000 samples to serve as the validation dataset.
Domain Dataset Name Size Reference
Code CodeFeedback 156k  Zheng et al. (2024b)
CodeAlpaca 20k Chaudhary (2023)
Math MathInstruction 262K Yue et al. (2023)
Medical MedMCQA 183k Zhang et al. (2023)
Law Legalbench 90k Guha et al. (2023)
Finance FinGPT 204k Wang et al. (2023b)
Alignment Ultrafeedback 62k Cui et al. (2024)
@@
Customization Training Hyperparameters. In model customization, we use different hyperparameters depending on the model size. For LLaMA2-70B, we apply QLoRA with the settings outlined in Table 11, while for 7B models, we use LoRA. For smaller models like Phi2 and Phi-1.5, we fine-tune all model parameters. For LLaMA2-70B, we fine-tune it as a quantized 4-bit model over 1 epoch, starting with a learning rate of $1.5\times 10^{-6}$. For the 7B models, we train for 3 epochs, with a seed value of 42. The training setup includes a weight decay of 0.1, a batch size of 128, a warmup ratio of 0.03, and input sequences of 512 tokens, following standard experimental practices (Hu et al., 2021). For Phi2 and Phi-1.5, we use the training hyperparameters from the LLaMA2 report. We employ the AdamW optimizer with a cosine learning rate scheduler, starting with a learning rate of $2\times 10^{-5}$, a weight decay of 0.1, a batch size of 128, and use bfloat16 precision for 512-token input sequences. Specifically, for alignment, we follow SimPO Meng et al. (2024) and set the preference parameters $\beta=2$ and $\gamma=1$. The learning rate is $1\times 10^{-6}$ for LLaMA2-70B and $5\times 10^{-7}$ for the 7B and smaller models. All experiments are conducted using the LLaMA-Factory on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, with PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
@@ B.8 Security Across Secure Sizes
50M $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v}$ $\mathrm{W}_{q},\mathrm{W}_{k},\mathrm{W}_{v},\mathrm{W}_{o}$ MLPMLP 0
@@
We further computed $\Delta\text{ADR}$ by close-sourcing varying quantities and proportions of parameters under FT-all attacks on three additional models. As shown in Figure 10 and Figure 11(a), we observed the same pattern as with Llama2-7B, where security emerges once a sufficient number of parameters are secured. For example, on Mistral-7B, security occurs after secure-sourcing 100 million parameters, which is less than a single decoder layer. Secure-sourcing fewer parameters leads to a notable drop in security, with $\Delta\text{ADR}$ rising to around 40%. Beyond this threshold, security stabilizes near 0% $\Delta\text{ADR}$. This pattern holds across all models, highlighting a critical threshold for effective secure-source. Furthermore, different architectures require varying secure-sourcing quantities to achieve security, even with similar model sizes. For instance, Mistral-7B reaches security by secure-sourcing 100 million parameters, Llama2-7B requires 200 million, and Phi-1.5 needs a higher rate of 7%, compared to 3% for Llama2-7B.
@@
We explore how secured parameter ratio impacts the model security in Llama2-7B, as shown in Figure 11(b). For instance, technical skills such as Math show earlier transitions, with security emerging at 1% parameters secured, whereas domains such as Commonsense Reasoning require hiding 3%. In summary, secure-sourcing a small portion of parameters can provide sufficient security against model distillation, meanwhile, technical capabilities tend to be more challenging to distill than other domains.
[…]